OpenStudy (anonymous):

Kool_kat: WriteTwo Equivalent Ratios For The Given Ratio 12/8

OpenStudy (katrinakaif):

12/8 = 3/2

OpenStudy (anonymous):

12/8 = 12*c/8*c ( choose any integer number c other than one like 2,3,

OpenStudy (anonymous):

?? raheen c ? im in 7th grade

OpenStudy (anonymous):

ok, I'm sorry 12/8 = (12÷ 2)/(8÷2) = 6/4 and so you can divid or multimly both the numerator and denomenator by or in the same number